OpenStudy (anonymous):

The speed of a passanger train is 12mph faster than the speed of a freight train. The passenger train travels 330 miles in the same time it takes the freight train to travel 270 miles. find the speed of each train?

OpenStudy (anonymous):

x+12=Y x/330=y/270 Now just solve the system :)

OpenStudy (anonymous):

330-270=60. so the trains travel fo 5 hours. 330/5=66 & 270/5=54. thats the answer! let's leave the jargon out of math please!

OpenStudy (anonymous):

jargon????

OpenStudy (anonymous):

stuff like "system of equations"... let's talk english as long as we can!

OpenStudy (anonymous):

Manoranjan, I do not apologize for using an appropriate level of discourse, and I do not appreciate your comments.

OpenStudy (anonymous):

sorry. no offence!

OpenStudy (anonymous):

apologies simplephysics!

OpenStudy (anonymous):

I really didnt mean to upset anyone over a math problem